OpenStudy (priyar):

what is the gain in potential energy of object of mass m raised from the surface of earth to height equal to radius R of the earth?

OpenStudy (priyar):

should i consider the variation in g with height???...because in some problems we consider and in some we don't...for ex:(here) at height=R,g'=g/4...so the answer will change....

OpenStudy (shamim):

Yes the value of gravitational acceleration g will b change

OpenStudy (shamim):

Can u calculate the work done for taking the object fron the surface of earth to the height R

OpenStudy (shamim):

U know Gain of potential energy=work done

OpenStudy (shamim):

U know g=GM/r^2

OpenStudy (shamim):

If r increase then the value of g will decrease

OpenStudy (priyar):

yes i know these..but i am getting the correct answer(as in the book) only if i consider that g wont change...

OpenStudy (priyar):

here is what i did.. g'=g/(1+h/R)^2=>g'=g/4 work done=-(final potential energy - initial potential energy) =-(-mg'R-0) =mg'R =mgR/4 =GMm/4R {since GM/R^2} but,when i do it in the below method i am getting the right answer work done=-(GMm/2R - GMm/R) {final P.E - initial P.E} =-(-GMm/2R) =GMm/2R what is wrong in the first method...??

OpenStudy (shamim):

U used g'=g/4

OpenStudy (shamim):

But when the height of ur object is h=R/2 Then g will b more than g/4

OpenStudy (shamim):

And the value of g will change when the object will go frm the surface of earth to the height h=R

OpenStudy (shamim):

So the value of g is a variable

OpenStudy (shamim):

U used g=g/4 is a constant

OpenStudy (shamim):

Do u know hw to calculate work done by a variable force

OpenStudy (priyar):

why height of the object is R/2?? @shamim

OpenStudy (priyar):

oh now i get it..g changes at each and every point...

OpenStudy (priyar):

yes i do know...W=integral(F.dr) W=integral 0toR (mgx/(1+x/R)^2.dx)...is this correct?

OpenStudy (farcher):

\[\Delta U = \dfrac{GM_Em}{R_E} - \dfrac{GM_Em}{2R_E}\]

OpenStudy (shamim):

Limit is correct

OpenStudy (shamim):

Limit will b 0 to R

OpenStudy (priyar):

ok what about the rest of the expression??

OpenStudy (priyar):

is the integral correct?

OpenStudy (priyar):

i tried this integration...but the answer does not math

OpenStudy (priyar):

So i will stick to the 2nd method....

OpenStudy (priyar):

@Farcher how to do this question with 'g'?

OpenStudy (shamim):

No u r not correct !!

OpenStudy (shamim):

U know Work done=gain in potential energy

OpenStudy (shamim):

Work done dw=F.dr

OpenStudy (shamim):

W=integration of Fdr

OpenStudy (priyar):

that is what i have done...right?

OpenStudy (shamim):

Limit is 0 to R

OpenStudy (shamim):

F=mg

OpenStudy (priyar):

i have written it from 0 to R only..

OpenStudy (shamim):

But g=GM/r^2

OpenStudy (priyar):

yes... i have substituted g as g/(1+x/R)^2 as it is constantly varying...

OpenStudy (shamim):

So W=integration of (GMm/r^2)dr

OpenStudy (priyar):

This formula comes in the chapter gravitation

OpenStudy (shamim):

Here u hv to use g=GM/r^2

OpenStudy (shamim):

U know G,M is constant

OpenStudy (shamim):

r is a variable

OpenStudy (shamim):

r is the distance between the centre of earth to the object

OpenStudy (priyar):

yes u are also correct...but i have done it using a formula which gives the variation of g with height..

OpenStudy (shamim):

So W=GMm* integration (r^-2)dr

OpenStudy (shamim):

W=GMm*(-1/r), limit 0 to R

OpenStudy (priyar):

but GMm*(1/0) is infinity??

OpenStudy (farcher):

To find the potential energy of a mass m at a distance r from a mass M you have to define a zero of potential . The zero of potential energy is usually taken as infinity. So the limits of integration from infinity to r. For this problem since you want a difference in potential energy you need to integrate from R_E to 2R_E. So there is no division by zero.

OpenStudy (shamim):

Sorry

OpenStudy (shamim):

Limit of integration will b R to 2R

OpenStudy (shamim):

Because When the object is on the surface of earth then r=R Because the distance frm the centre of earth to the object is =R

OpenStudy (shamim):

When the object is at height R Then r=2R Because the distance frm the centre of earth to the object is=2R

OpenStudy (priyar):

thanks @Farcher and @shamim ! but haven't u heard of this formula: g'=g/(1+h/R)^2 for finding gravitational acceleration at ht. h ?

OpenStudy (shamim):

I know ur formula

OpenStudy (shamim):

In ur formula h is variable

OpenStudy (shamim):

Nd h is the distance frm the surface of earth to the object

OpenStudy (shamim):

We need the distance frm the centre of earth to the object

OpenStudy (shamim):

So we need r

OpenStudy (shamim):

not h

OpenStudy (priyar):

ok..